Hi! Hoping you could help me solve this one. Thank you in advance! :)
An agent receives a weekly salary of P3,200 plus 15% of his total sales. If for a given week he got P5,450, how much was his total sales?
This is the formula I found in my book (although I have it, I still don't know how to solve it):
Commisson (C) = Rate (R) x Total Sales
Thank you! :)

Answer by stanbon(75887)

An agent receives a weekly salary of P3,200 plus 15% of his total sales. If for a given week he got P5,450, how much was his total sales?
This is the formula I found in my book (although I have it, I still don't know how to solve it):
Commission (C) = Rate (R) x Total Sales
-----
Total Earnings = 3200 + 0.15*sales
5450 = 3200 + 0.15*S
2350 = 0.15*S
S = 2350/0.15
Ans: Total sales = $15666.67
-----------
Cheers,
Stan H.
--------